# Solving Harder Reinforcement Learning Problems

### v3

@@ 1 Introduction
Recent years have seen the rapid rise of large language models (LLMs) trained with internet-scale pretraining data [33] with post training using both supervised fine-tuning [40] and reinforcement learning (RL) [1, 37, 3, 18]. The seminal paradigm of RL with human feedback [9] is limited by the human-derived data it is based on and the reward hacking issues that arise from the use of subjective signals more generally [2]. To enable progress toward superhuman capabilities, current work is focusing on grounded reward signals that are free of fine-grained human input as in code generation [35, 28, 15, 45, 19] and mathematics [25, 5, 13, 44].
The policy gradients family of RL methods [41] has proven effective in language model training [18]. To scale to new capabilities, model training needs to tackle challenging RL task sets with no known solutions but for which correctness may be verified, as in formal mathematics environments [5]. In such settings, the RL training loop both updates model parameters and searches for solutions to problems at the continuously advancing frontier of model capabilities.
The specific search algorithm introduces a coupling between inference and model updates, which means that naively optimizing the expected single sample reward, or $\mathrm{pass@1}$ may be sub-optimal. While various inference-time search methods are possible [22, 24, 27, 42], simply taking multiple independent samples from the model has proven rather effective [30]. Our contribution is to couple this simple search method with model parameter updates by enabling robust optimization of the $\mathrm{pass@k}$ objective, which is the expected maximum reward over $k$ independent samples.
@@
The $\mathrm{pass@k}$ was championed by [11] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1, generalize to continuous rewards in Theorem 3, and provide additional characterisations in Corollaries 2 and 3. Concurrently with our work [39] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting $\mathrm{pass@k}$ in terms of a partial sort, [14] and [43] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-$N$ [10, 6] and other more general [7] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-$N$ prediction with a single sample was addressed by [34]. Our contribution can be interpreted as a generalization and variance reduction of [39]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [29, 31].
@@
Our theme is constructing robust estimators of the $\mathrm{pass@k}$ and its gradient given $n\geq k$ samples by averaging (over all $\binom{n}{k}$ subsets of size $k$) simple estimators that are functions of $k$ samples. This is straightforward for binary rewards (Section 2), using the counting proof of Theorem 4. We generalize to continuous rewards using the key trick of assuming without loss of generality that the rewards are sorted, as in Section 3. Finally, we give baselining methods that require more involved derivations due to averaging over all subsets that do not include a given element (to retain unbiasedness) but which boil down to the same easy-to-apply results in Section 4, yielding our Pass-at-$k$ Policy Optimization (PKPO). We present toy experiments in Section 5.1 which demonstrate the variance reduction afforded by our estimators. Finally, Section 5.2 demonstrates that using our reward transformation solves more tasks and selectively optimizes $\mathrm{pass@k}$ through RL experiments on Gemma2 [38] and Llama3.1 [16], showcasing real-world impact.
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
An estimator for the $\mathrm{pass@k}$ was given in [11]: given $n\geq k$ i.i.d. samples of which $c$ are correct, the estimator is
@@
The following was proven in [11]; we give a different proof that sets up our gradient estimator.
@@ 3 Continuous Rewards
To compute this stably we cancel factors in the binomial coefficients to get
$\rho^{(g)}(n,c,k)\equiv\frac{k}{n-k+1}\sum_{i=k}^{n}g_{i}\prod_{j=1}^{k-1}\frac{i-j+1}{n-j+1}.$ (14)
@@ 4.1 Leave-One-Out Baseline for the Simple Case
A popular variance reduction method [29, 31, 18] for point-wise rewards $g(x)$ subtracts the mean of the leave one out (LOO) rewards within each mini-batch $x_{1},x_{2},\dots,x_{n}$:
@@
By assuming w.l.o.g. an ascending ordering of $g(x_{i})$, excluding the first index does not change the ordering of the remaining indices. The first term is therefore
$b_{1}^{(k)}=\sum_{i=2}^{N}s_{i}=\frac{1}{\binom{n}{k}}\sum_{i=2}^{N}\big(m_{ii}+(i-2)m_{i,i+1}\big)g(x_{i}),$ (31)
@@
The baseline $b_{i}^{(k)}$ is an average of the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ estimates over sets of size $k$. For a number of samples $n$ equal to $k$, there are no such subsets to construct the baseline. [39] recently overcame this issue for the specific case $n=k$ by using $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@(k-1)}$ as the baseline statistic. We generalize their approach to $k<n$ and to averaging over all subsets by defining similarly to Equation 27
@@ 5.2 RL on Open Source LLMs
We demonstrate promising RL results with the 2B and 9B parameter variants of Gemma2 [38] and the 8B parameter variant of Llama3.1 on real-world problems in MATH [20], code generation [4] [12], and the easy public subset of ARC-AGI-1 [8]. The latter is a challenging reasoning task-set even for state-of-the-art models much larger than Gemma2.
For Gemma2-2B we use a v5litepod-128 [17] which needs around 4 hours per 1000 training steps. Each RL training run [36] involves sampling a fixed $n$ number of completions $\{x_{i}\}_{i=1}^{n}$ for a given prompt at a given training step. For our experiments, we set $n=16$. The rewards are computed for every completion using a reward function $g(\cdot)$. We transform these rewards $\{g(x_{i})\}_{i=1}^{n}$ using our unbiased estimator $s^{(\mathrm{loo}-1)}$ of (33), which we favour due to Figure 4, and which we refer to as PKPO. We repeat the training for a selection of $k^{\mathrm{opt}}$, thus optimizing a different $\mathrm{pass@k^{opt}}$ each time. Since $k^{\mathrm{opt}}=1$ leads to no reward transformation, this is our baseline (although we use basic LOO mean centering of Equation 26, without which the training diverges). For each run, we measure $\mathrm{pass@k^{eval}}$ for every $k^{\mathrm{eval}}$$\in\{1,2,4,8,12,16\}$ at each step. Additionally, we also track model entropy and cumulative solve rate during training. The latter is defined as the fraction of tasks from the task-set for which the model has sampled a correct solution at least once; this is a critical metric that reflects the success of the model’s exploration and measures its ability to find novel solutions.
@@
Entropy regularization baseline In addition to our PKPO and the special case thereof of [39], we also add the entropy regularization baseline, which is PPO with an additional entropy term in the objective. We give this baseline an arguably unfair advantage by performing a small sweep over the values $0.001,0.005,0.01,0.05,0.1$ for the entropy_coefficient for each (model, benchmark) pair and only report the best result as EntropyReg.
@@ 5.2.1 Choosing $k^{\mathrm{opt}}$ selectively optimizes $\mathrm{pass@k^{eval}}$ and solves more tasks
We use the training split of Hendrycks MATH [20] which contains 12,000 problems as our task set. Figure 6(a) shows that a higher $k^{\mathrm{opt}}$ in our transformation leads to a consistently higher cumulative solve rate throughout training, as well as a higher entropy. By optimizing $\mathrm{pass@k}$ instead of $\mathrm{pass@1}$, the model appears to better utilize the exploration budget thus finding more solutions.
In Figure 7, we compare $\mathrm{pass@k^{eval}}$ across our runs ($k^{\mathrm{opt}}$ $\in\{1,4,8\}$) for various $k^{\mathrm{eval}}$. We find the best $\mathrm{pass@k^{eval}}$ when $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ (or $k^{\mathrm{opt}}$ is closest to $k^{\mathrm{eval}}$ among available $k^{\mathrm{opt}}$). Non-transformed rewards optimize $\mathrm{pass@1}$, leading to sub-optimal $\mathrm{pass@k^{eval}}$ for $k^{\mathrm{eval}}$$\neq 1$, and the deficit worsens as $k^{\mathrm{eval}}$ increases. Thus, our experiments also demonstrate that setting $k^{\mathrm{opt}}\coloneqq$ $k^{\mathrm{eval}}$ in our transformation suffices to optimize $\mathrm{pass@k^{eval}}$ for a $k^{\mathrm{eval}}$$\leq n$. This generalizes the already powerful result of [39] by alleviating the coupling that restricts to optimizing either $\mathrm{pass@n}$ or $\mathrm{pass@1}$. In other words, since RL training of LLMs typically samples a large batch ($n\gg 1$), failing to use our transformation results in sub-optimal $\mathrm{pass@k}$ performance, especially for modest values of $k$.
As $k^{\mathrm{opt}}\longrightarrow n$, the variance of our estimator increases as there are fewer subsets in (33) (see Figure 4). We presume this is why 1) gains of $k^{\mathrm{opt}}=8$ over $k^{\mathrm{opt}}=4$ are more prominent when $k^{\mathrm{eval}}$ $\in\{12,16\}$ than when $k^{\mathrm{eval}}$ =8=8. That is, when $k^{\mathrm{eval}}$ is further away from $k^{\mathrm{opt}}\in\{4,8\}$ than when it is closer, and 2) the special case $n=k^{\mathrm{opt}}$ of [39] struggles to optimize the $\mathrm{pass@n}$.
@@ 5.2.2 PKPO robustly improves $\mathrm{pass@k}$ on held out evaluations
Tables 1-4 above (and Tables 7-8 in the appendix) present performance on held-out sets for two tasks. We report the mean and standard error based on three runs with different random seeds. For math, we train on the train split and evaluate on the test split of Hendrycks MATH [20]. To evaluate coding, we use MBPP [4] for training and evaluate on HumanEval [12]. MBPP has multiple unit tests per problem and hence we use this not only as a proxy for additional benchmarks but also to showcase our handling of a continuous reward function (% unit tests passed).
@@ 5.2.3 Improving $\mathrm{pass@k}$ without sacrificing $\mathrm{pass}@1$
Figure 3 demonstrates that as PKPO can use any arbitrary $k^{\mathrm{opt}}\leq n$, this allows varying $k^{\mathrm{opt}}$ over the course of training to good effect. We show a simple annealing procedure which starts training with a high $k^{\mathrm{opt}}=8$ and reduces it to $k^{\mathrm{opt}}=1$ after 1500 steps. This trains the model to initially prioritize exploration (optimize $\mathrm{pass@k}$) and then consolidate the single-sample policy (optimize $\mathrm{pass@1}$). This switch is apparent in Figure 3(a), at step 1500 where the slope of $k^{\mathrm{annealed}}$ changes. While traditional methods like [39] suffer from a trade-off between $\mathrm{pass@k}$ and $\mathrm{pass@1}$, we get a final model which has higher $\mathrm{pass@k^{eval}}$ for all $k^{\mathrm{eval}}$>1>1 with no sacrifice in $\mathrm{pass@1}$.
@@ 5.2.4 PKPO is essential for learning on hard problems
Figure 8 shows the limitation of traditional $\mathrm{pass@1}$ optimization through RL on an especially challenging task-set. We use the easy subset of ARC-AGI-1 [8]. We observe that conventional $\mathrm{pass@1}$ optimization stalls. However, our $\mathrm{pass@k}$ approach unblocks learning, and results in higher $\mathrm{pass@k^{eval}}$ across all $k^{\mathrm{eval}}$ including $k^{\mathrm{eval}}$=1=1. Furthermore, we see higher $k^{\mathrm{opt}}$ leads to more effective and faster learning. This is likely because the benefits of prioritizing joint utility over individual sample utility are more prominent on a harder task-set.
@@
Tables 5 and 6 show more extensive experiments on ARC-AGI-1. We make an 80:20 train:test split of the same easy subset as before and report the cumulative solve rate on the train set and $\mathrm{pass@k}$ rate on the test set. We train to saturation (no change in cumulative rate for 1k steps), and again use three random restarts to provide standard errors. By encouraging exploration in a direct and stable manner, our method unblocks learning unlike other methods. Entropy Regularization does indeed sacrifice pass@1 and slightly improves $\mathrm{pass@k}$ by promoting exploration, but it is hard to tune, and is significantly outperformed by our method. Moreover, it has no explicit way to optimize for a specific k_eval. [39] targets the same objective as PKPO, but couples the minibatch size to $k$ and thereby incurs higher variance than PKPO with $k<n$.

### v4

@@ 1 Introduction
Recent years have seen the rapid rise of large language models (LLMs) trained with internet-scale pretraining data [pretraining] with post training using both supervised fine-tuning [posttrainingsft] and reinforcement learning (RL) [gpt4tr, gemini2023, claude, deepseekr1]. The seminal paradigm of RL with human feedback [rlhf] is limited by the human-derived data it is based on and the reward hacking issues that arise from the use of subjective signals more generally [rewardhacking]. To enable progress toward superhuman capabilities, current work is focusing on grounded reward signals that are free of fine-grained human input as in code generation [shojaee2023execution, le2022coderl, dou2024stepcoder, yu2023b, gehring2024rlef] and mathematics [alphacode, alphaproofblog, alphageometry2, yang2023leandojo].
The policy gradients family of RL methods [policygradients] has proven effective in language model training [deepseekr1]. To scale to new capabilities, model training needs to tackle challenging RL task sets with no known solutions but for which correctness may be verified, as in formal mathematics environments [alphaproofblog]. In such settings, the RL training loop both updates model parameters and searches for solutions to problems at the continuously advancing frontier of model capabilities.
The specific search algorithm introduces a coupling between inference and model updates, which means that naively optimizing the expected single sample reward, or $\mathrm{pass@1}$ may be sub-optimal. While various inference-time search methods are possible [inftimea, inftimeb, inftimec, inftimed], simply taking multiple independent samples from the model has proven rather effective [iidbeatsselfrepair]. Our contribution is to couple this simple search method with model parameter updates by enabling robust optimization of the $\mathrm{pass@k}$ objective, which is the expected maximum reward over $k$ independent samples.
@@
The $\mathrm{pass@k}$ was championed by [chen2021codex] who gave a popular unbiased estimator of the metric which we derive from a new perspective (to set up our gradient estimators) as Theorem 1, generalize to continuous rewards in Theorem 3, and provide additional characterisations in Corollaries 2 and 3. Concurrently with our work [tangloo] offered an elegant variance reduction method for the gradient of the $\mathrm{pass@k}$ which corresponds to the special case $n=k$ of our Equation 33. Interpreting $\mathrm{pass@k}$ in terms of a partial sort, [differentiable_ranking_and_sorting_with_ot] and [differentiable_ranking_and_sorting_with_ot_xie] present elegant approximations that are rather general but less efficient in our setting. Others have provided variational approximations for handling the closely related Best-of-$N$ [bonaware, variationalbon] and other more general [infalign] inference-time algorithms. The contrasting idea of training a model to approximate the Best-of-$N$ prediction with a single sample was addressed by [bond]. Our contribution can be interpreted as a generalization and variance reduction of [tangloo]. For a general discussion of gradient estimation, variance reduction, and Monte Carlo, we recommend [shakirreview, mcbook].
@@
Our theme is constructing robust estimators of the $\mathrm{pass@k}$ and its gradient given $n\geq k$ samples by averaging (over all $\binom{n}{k}$ subsets of size $k$) simple estimators that are functions of $k$ samples. This is straightforward for binary rewards (Section 2), using the counting proof of Theorem 4. We generalize to continuous rewards using the key trick of assuming without loss of generality that the rewards are sorted, as in Section 3. Finally, we give baselining methods that require more involved derivations due to averaging over all subsets that do not include a given element (to retain unbiasedness) but which boil down to the same easy-to-apply results in Section 4, yielding our Pass-at-$k$ Policy Optimization (PKPO). We present toy experiments in Section 5.1 which demonstrate the variance reduction afforded by our estimators. Finally, Section 5.2 demonstrates that using our reward transformation solves more tasks and selectively optimizes $\mathrm{pass@k}$ through RL experiments on Gemma2 [gemma2] and Llama3.1 [llama], showcasing real-world impact.
@@ 2.1 An Unbiased $\mathrm{pass@k}$ Estimator
An estimator for the $\mathrm{pass@k}$ was given in [chen2021codex]: given $n\geq k$ i.i.d. samples of which $c$ are correct, the estimator is
@@
The following was proven in [chen2021codex]; we give a different proof that sets up our gradient estimator.
@@ 3 Continuous Rewards
To compute this stably we cancel factors in the binomial coefficients to get [^2] We thank to Ruixu Zhou of Tsinghua University for correcting errors in equations 14, 31 and 32.
$\rho^{(g)}(n,c,k)\equiv\frac{k}{n-k+1}\sum_{i=k}^{n}g_{i}\prod_{j=1}^{k-1}\frac{i-j}{n-j+1}.$ (14)
@@ 4.1 Leave-One-Out Baseline for the Simple Case
A popular variance reduction method [shakirreview, mcbook, deepseekr1] for point-wise rewards $g(x)$ subtracts the mean of the leave one out (LOO) rewards within each mini-batch $x_{1},x_{2},\dots,x_{n}$:
@@
[…]
$b_{1}^{(k)}=\sum_{i=2}^{N}s_{i}=\frac{1}{\binom{n-1}{k}}\sum_{i=2}^{N}\big(m_{ii}+m_{i-1,i}(i-2)\big)g(x_{i}),$ (31)
@@
The baseline $b_{i}^{(k)}$ is an average of the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ estimates over sets of size $k$. For a number of samples $n$ equal to $k$, there are no such subsets to construct the baseline. [tangloo] recently overcame this issue for the specific case $n=k$ by using $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@(k-1)}$ as the baseline statistic. We generalize their approach to $k<n$ and to averaging over all subsets by defining similarly to Equation 27
@@ 5.2 RL on Open Source LLMs
We demonstrate promising RL results with the 2B and 9B parameter variants of Gemma2 [gemma2] and the 8B parameter variant of Llama3.1 on real-world problems in MATH [math_12000], code generation [austin2021programsynthesislargelanguage] [chen2021evaluatinglargelanguagemodels], and the easy public subset of ARC-AGI-1 [arc_agi_1]. The latter is a challenging reasoning task-set even for state-of-the-art models much larger than Gemma2.
For Gemma2-2B we use a v5litepod-128 [gcp] which needs around 4 hours per 1000 training steps. Each RL training run [ppo_rl] involves sampling a fixed $n$ number of completions $\{x_{i}\}_{i=1}^{n}$ for a given prompt at a given training step. For our experiments, we set $n=16$. The rewards are computed for every completion using a reward function $g(\cdot)$. We transform these rewards $\{g(x_{i})\}_{i=1}^{n}$ using our unbiased estimator $s^{(\mathrm{loo}-1)}$ of (33), which we favour due to Figure 4, and which we refer to as PKPO. We repeat the training for a selection of $k^{\mathrm{opt}}$, thus optimizing a different $\mathrm{pass@k^{opt}}$ each time. Since $k^{\mathrm{opt}}=1$ leads to no reward transformation, this is our baseline (although we use basic LOO mean centering of Equation 26, without which the training diverges). For each run, we measure $\mathrm{pass@k^{eval}}$ for every $k^{\mathrm{eval}}$$\in\{1,2,4,8,12,16\}$ at each step. Additionally, we also track model entropy and cumulative solve rate during training. The latter is defined as the fraction of tasks from the task-set for which the model has sampled a correct solution at least once; this is a critical metric that reflects the success of the model’s exploration and measures its ability to find novel solutions.
@@
Entropy regularization baseline In addition to our PKPO and the special case thereof of [tangloo], we also add the entropy regularization baseline, which is PPO with an additional entropy term in the objective. We give this baseline an arguably unfair advantage by performing a small sweep over the values $0.001,0.005,0.01,0.05,0.1$ for the entropy_coefficient for each (model, benchmark) pair and only report the best result as EntropyReg.
@@ 5.2.1 Choosing $k^{\mathrm{opt}}$ selectively optimizes $\mathrm{pass@k^{eval}}$ and solves more tasks
We use the training split of Hendrycks MATH [math_12000] which contains 12,000 problems as our task set. Figure 6(a) shows that a higher $k^{\mathrm{opt}}$ in our transformation leads to a consistently higher cumulative solve rate throughout training, as well as a higher entropy. By optimizing $\mathrm{pass@k}$ instead of $\mathrm{pass@1}$, the model appears to better utilize the exploration budget thus finding more solutions.
In Figure 7, we compare $\mathrm{pass@k^{eval}}$ across our runs ($k^{\mathrm{opt}}$ $\in\{1,4,8\}$) for various $k^{\mathrm{eval}}$. We find the best $\mathrm{pass@k^{eval}}$ when $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ (or $k^{\mathrm{opt}}$ is closest to $k^{\mathrm{eval}}$ among available $k^{\mathrm{opt}}$). Non-transformed rewards optimize $\mathrm{pass@1}$, leading to sub-optimal $\mathrm{pass@k^{eval}}$ for $k^{\mathrm{eval}}$$\neq 1$, and the deficit worsens as $k^{\mathrm{eval}}$ increases. Thus, our experiments also demonstrate that setting $k^{\mathrm{opt}}\coloneqq$ $k^{\mathrm{eval}}$ in our transformation suffices to optimize $\mathrm{pass@k^{eval}}$ for a $k^{\mathrm{eval}}$$\leq n$. This generalizes the already powerful result of [tangloo] by alleviating the coupling that restricts to optimizing either $\mathrm{pass@n}$ or $\mathrm{pass@1}$. In other words, since RL training of LLMs typically samples a large batch ($n\gg 1$), failing to use our transformation results in sub-optimal $\mathrm{pass@k}$ performance, especially for modest values of $k$.
As $k^{\mathrm{opt}}\longrightarrow n$, the variance of our estimator increases as there are fewer subsets in (33) (see Figure 4). We presume this is why 1) gains of $k^{\mathrm{opt}}=8$ over $k^{\mathrm{opt}}=4$ are more prominent when $k^{\mathrm{eval}}$ $\in\{12,16\}$ than when $k^{\mathrm{eval}}$ =8=8. That is, when $k^{\mathrm{eval}}$ is further away from $k^{\mathrm{opt}}\in\{4,8\}$ than when it is closer, and 2) the special case $n=k^{\mathrm{opt}}$ of [tangloo] struggles to optimize the $\mathrm{pass@n}$.
@@ 5.2.2 PKPO robustly improves $\mathrm{pass@k}$ on held out evaluations
Tables 1-4 above (and Tables 7-8 in the appendix) present performance on held-out sets for two tasks. We report the mean and standard error based on three runs with different random seeds. For math, we train on the train split and evaluate on the test split of Hendrycks MATH [math_12000]. To evaluate coding, we use MBPP [austin2021programsynthesislargelanguage] for training and evaluate on HumanEval [chen2021evaluatinglargelanguagemodels]. MBPP has multiple unit tests per problem and hence we use this not only as a proxy for additional benchmarks but also to showcase our handling of a continuous reward function (% unit tests passed).
@@ 5.2.3 Improving $\mathrm{pass@k}$ without sacrificing $\mathrm{pass}@1$
Figure 3 demonstrates that as PKPO can use any arbitrary $k^{\mathrm{opt}}\leq n$, this allows varying $k^{\mathrm{opt}}$ over the course of training to good effect. We show a simple annealing procedure which starts training with a high $k^{\mathrm{opt}}=8$ and reduces it to $k^{\mathrm{opt}}=1$ after 1500 steps. This trains the model to initially prioritize exploration (optimize $\mathrm{pass@k}$) and then consolidate the single-sample policy (optimize $\mathrm{pass@1}$). This switch is apparent in Figure 3(a), at step 1500 where the slope of $k^{\mathrm{annealed}}$ changes. While traditional methods like [tangloo] suffer from a trade-off between $\mathrm{pass@k}$ and $\mathrm{pass@1}$, we get a final model which has higher $\mathrm{pass@k^{eval}}$ for all $k^{\mathrm{eval}}$>1>1 with no sacrifice in $\mathrm{pass@1}$.
@@ 5.2.4 PKPO is essential for learning on hard problems
Figure 8 shows the limitation of traditional $\mathrm{pass@1}$ optimization through RL on an especially challenging task-set. We use the easy subset of ARC-AGI-1 [arc_agi_1]. We observe that conventional $\mathrm{pass@1}$ optimization stalls. However, our $\mathrm{pass@k}$ approach unblocks learning, and results in higher $\mathrm{pass@k^{eval}}$ across all $k^{\mathrm{eval}}$ including $k^{\mathrm{eval}}$=1=1. Furthermore, we see higher $k^{\mathrm{opt}}$ leads to more effective and faster learning. This is likely because the benefits of prioritizing joint utility over individual sample utility are more prominent on a harder task-set.
@@
Tables 5 and 6 show more extensive experiments on ARC-AGI-1. We make an 80:20 train:test split of the same easy subset as before and report the cumulative solve rate on the train set and $\mathrm{pass@k}$ rate on the test set. We train to saturation (no change in cumulative rate for 1k steps), and again use three random restarts to provide standard errors. By encouraging exploration in a direct and stable manner, our method unblocks learning unlike other methods. Entropy Regularization does indeed sacrifice pass@1 and slightly improves $\mathrm{pass@k}$ by promoting exploration, but it is hard to tune, and is significantly outperformed by our method. Moreover, it has no explicit way to optimize for a specific k_eval. [tangloo] targets the same objective as PKPO, but couples the minibatch size to $k$ and thereby incurs higher variance than PKPO with $k<n$.